OpenStudy (anonymous):

solve the equation 7x^2 + 6 = 13 ?

OpenStudy (uri):

Subract 6 both sides 7x^2 + 6 = 13 -6 -6 ========== What do you get? @broderick365

OpenStudy (uri):

There...

OpenStudy (anonymous):

-7

OpenStudy (uri):

Why -7?

OpenStudy (uri):

13 the greater number has positive sign.. So, 7x^2=7

OpenStudy (uri):

Now divide 7 with 7,what do you get?

OpenStudy (anonymous):

1

OpenStudy (uri):

Yes now the answer will be x=+1 -1